Question 7Entries for Issuing No-Par Stock On May 15, Helena Carpet Inc., a carpet wholesaler, issued for cash 305,000 shares of no-par
Answer #1
a
May 15 Cash 3660000 =305000*12
     Common Stock 1220000 =305000*4
     Paid in Capital in Excess of stated value-Common Stock 2440000
June 30 Cash 1740000 =20000*87
     Preferred Stock 1600000 =20000*80
     Paid in Capital in Excess of Par-Preferred Stock 140000
b
Total paid-in Capital 5400000 =3660000+1740000